OpenStudy (anonymous):

i have (1,5) and (3,1) as my given points how can i get my third point with only the given information....and how could i get my slope

OpenStudy (anonymous):

You have to find the equation of the line,for that you need to find the slope first which is -2 here,equation is \[ y-5=-2(x-1) \] \[ 2x +y = 7\]